Differentiate the following functions. Show step by step solution and

Indicate your final answer.

1. X = 1/t

2. y = 4x2– 3x – 2

3. y = (x+2) 1/2

1
Expert's answer
2022-03-08T00:55:03-0500

Let us differentiate the following functions.


1. "X '=(\frac{ 1}t)'=-\frac{1}{t^2}"


2. "y '= (4x^2\u2013 3x \u2013 2)'=8x-3"


3. "y' = ((x+2) ^\frac{1}2)'=\frac{1}2(x+2)^{-\frac{1}2}"